OpenStudy (anonymous):

PLEASE HELP ME Of 357 randomly selected medical students, 30 said that they planned to work in a rural 1) community. Find a 95% confidence interval for the true proportion of all medical students who plan to work in a rural community.

OpenStudy (kropot72):

The 95% confidence interval for the true proportion is found from the following formula: \[p-1.96\sqrt{\frac{p(1-p)}{n}} < \pi < p+1.96\sqrt{\frac{p(1-p)}{n}}\] where p =357 and n = 357

OpenStudy (anonymous):

so i just fill in the numbers and that will be my answer??

OpenStudy (kropot72):

Basically yes. You will need to do some calculations to get the two values of probability that are the upper and lower limit.

OpenStudy (anonymous):

and both p and n are 357?

OpenStudy (kropot72):

Sorry, my bad. p = 30/357 n = 357

OpenStudy (anonymous):

30 divided 357?

OpenStudy (kropot72):

\[p=\frac{30}{357}\] \[n=357\]